### v1

# A general frame of quantum simulation for nonlinear partial differential equations

Abstract Currently, Jin et al. proposed a quantum simulation technique for any a linear PDE, called Schrödingerisation [1, 2, 3], which has been successfully applied to solve many non-Hamiltonian linear PDEs. In this paper, the Schrödingerisation technique of quantum simulation is expanded to any a nonlinear PDE by means of combining the Schrödingerisation technique with the homotopy analysis method (HAM) [4, 5, 6] that can transfer any a nonlinear PDE into a series of linear PDEs with convergence guarantee of series solution. In this way, a nonlinear PDE can be solved by quantum simulation using a quantum computer – yet to be developed in the future. For simplicity, we call it “the HAM-Schrödingerisation quantum algorithm”.
1 Introduction
Today quantum simulation [7, 8, 9, 10, 11, 12, 13] is a central issue in computational methods. Hamiltonian simulation technique might be most important in quantum computing algorithms, which is valid for a time-dependent Schrödinger equation
$i\;\partial_{t}\psi=H(t)\;\psi,$ (1)
where $H(t)$ is a time-dependent Hamiltonian operator and $i=\sqrt{-1}$.
Currently, a generic framework coined as Schrödingerisation [1, 2, 3] was introduced by Shi Jin et al., which can map any linear partial differential equations (PDEs) into Schrödinger equations in real time. This is a milestone in quantum simulation. Based on a new transformation called “warped phase transformation”, Schrödingerisation can solve any linear PDEs with quantum simulation in a rather general form
$\frac{\partial\psi({\bf r},t)}{\partial t}={\cal L}[\psi({\bf r},t)]+f({\bf r},t),\;\;\;{\bf r}\in\Omega,t\geq 0,$ (2)
subject to the initial condition
$\psi({\bf r},0)=\alpha({\bf r})$ (3)
and the boundary condition
$\psi({\bf r},t)=\beta(t)\;\;\mbox{when ${\bf r}\in\Gamma$},$ (4)
where $\bf r$ and $t$ denote the spatial and temporal variable, $\psi({\bf r},t)$ is a unknown function, $\cal L$ is a linear operator, $f({\bf r},t),\alpha({\bf r})$ and $\beta(t)$ are known functions, $\Omega$ and $\Gamma$ denote the physical domain and boundary, respectively. In practice, a linear PDE in the form (2) - (4) can be discretized in space to get a system of linear ordinary differential equations (ODEs) as follows:
$\frac{d{\bf u}(t)}{dt}=A(t){\bf u}(t)+{\bf b}(t),\;\;{\bf u}(0)={\bf a},$ (5)
where ${\bf b}\in\mathbb{C}^{n}$ and ${\bf u}\in\mathbb{C}^{n}$ are known functions, ${\bf a}\in\mathbb{C}^{n}$ is a known vector, $A(t)\in\mathbb{C}^{n\times n}$ can be a non-Hermitian matrix, i.e. $A(t)$ might be not equal to its conjugate transpose, respectively. The key point is that the linear PDEs (2)-(4) in the rather general form can be solved by the quantum simulation technique coined as Schrödingerisation [1, 2, 3]. Some applications illustrate the validity of Schrödingerisation for many types of linear PDEs [14, 15, 16, 17, 18, 19, 20, 21, 22, 23].
Can the Schrödingerisation technique of quantum simulation for any linear PDEs [1, 2, 3] be expanded to any nonlinear PDEs? The answer is positive, as briefly described below.
2 HAM-based quantum simulation for nonlinear PDEs
Can any a nonlinear PDE be transferred into a series of linear PDEs with convergence guarantee of solution series? The answer is positive, as described below in details.
In 1992 Liao [4] proposed a new analytic approximation method for highly nonlinear equations (including ODEs, PDEs and all other types), namely the homotopy analysis method (HAM) [5, 6, 24, 25, 26, 27, 28, 29, 30]. Based on homotopy, a basic concept in topology that describes continuous deformations, the HAM can transfer any a nonlinear equation into a series of linear sub-equations without any physical assumptions: the solution of the original nonlinear equations is equal to the summation of the solutions of these linear sub-equations. Especially, unlike perturbation methods and all other approximation techniques for nonlinear equations, the convergence of the solution series given by the HAM is guaranteed by means of an artificial parameter, called “convergence-control parameter” $c_{0}$. Besides, unlike perturbation methods, the HAM provides the great freedom to choose the type of the linear sub-equations [31, 32, 30] and also the great freedom to choose an initial guess solution. Up to now, several thousands of articles related to the HAM have been published in a wide range of fields including applied mathematics, physics, engineerings, nonlinear mechanics, quantum mechanics, bio-mechanics, astronomy, finances and so on [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46].
Without loss of generality, let us consider here a general nonlinear PDE
$\frac{\partial\psi({\bf r},t)}{\partial t}={\cal N}[\psi({\bf r},t)]+g({\bf r},t),\;\;{\bf r}\in\Omega,t\geq 0,$ (6)
subject to the initial condition
$\psi({\bf r},0)=\alpha({\bf r})$ (7)
and the boundary condition
$\psi({\bf r},t)=\beta(t)\;\;\mbox{when ${\bf r}\in\Gamma$},$ (8)
where $\bf r$ and $t$ denote the spatial and temporal variable, $\psi({\bf r},t)$ is a unknown function, $\cal N$ is a nonlinear operator, $g({\bf r},t),\alpha({\bf r})$ and $\beta(t)$ are known functions, $\Omega$ and $\Gamma$ denote the physical domain and boundary, respectively.
In the frame of the HAM, let $q\in[0,1]$ denote the embedding parameter that has no physical meanings, $\psi_{0}({\bf r},t)$ be a guess solution of $\psi({\bf r},t)$, the non-zero constant $c_{0}$ be the so-called “convergence-control parameter”, ${\mathcal{L}}^{\star}$ denote an auxiliary linear operator with the property ${\mathcal{L}}^{\star}[0]=0$, respectively. It should be emphasized here that one has great freedom to choose the guess solution $\psi_{0}({\bf r},t)$ and the auxiliary linear operator ${\cal L}^{\star}$. Then, a family of solution $\Psi({\bf r},t,q)$, where $q\in[0,1]$ is an embedding parameter in topology, can be constructed by the so-called “zeroth-order deformation equation”
$(1-q){\cal L}^{\star}\left[\Psi({\bf r},t,q)-\psi_{0}({\bf r},t)\right]=c_{0}\;q\;\left\{\frac{\partial\psi({\bf r},t)}{\partial t}-{\cal N}[\psi({\bf r},t)]-g({\bf r},t)\right\},$ (9)
subject to the initial condition
[…]
and the boundary condition
$\Psi({\bf r},t,q)=(1-q)\;\psi_{0}({\bf r},t)+q\;\beta(t)\hskip 21.33955pt\mbox{when ${\bf r}\in\Gamma$}.$ (11)
Due to the property ${\cal L}^{\star}[0]=0$, the solution of Eqs. (9) - (11) when $q=0$ is exactly the initial guess, i.e.
$\Psi({\bf r},t,0)=\psi_{0}({\bf r},t).$ (12)
When $q=1$, Eqs. (9) - (11) are exactly the same as the original nonlinear PDEs (6)-(8), thus
$\Psi({\bf r},t,1)=\psi({\bf r},t).$ (13)
So, the solution $\Psi({\bf r},t,q)$ of Eqs. (9) - (11), where $q\in[0,1]$ is the embedding parameter, connects the known guess solution $\psi_{0}({\bf r},t)$ and the unknown solution $\psi({\bf r},t)$ of the original nonlinear PDEs (6)-(8). Since one has great freedom [31, 32, 30] to choose the guess solution $\psi_{0}({\bf r},t)$ and the auxiliary linear operator ${\cal L}^{\star}$, it is reasonable to assume that both of $\psi_{0}({\bf r},t)$ and ${\cal L}^{\star}$ are so properly chosen that $\Psi({\bf r},t,q)$ is a continuous deformation with respect to the embedding parameter $q\in[0,1]$, say, as $q$ increases from 0 to 1, $\Psi({\bf r},t,q)$ deforms continuously from the known guess solution $\psi_{0}({\bf r},t)$ to the unknown solution $\psi({\bf r},t)$ of the original nonlinear PDEs (6)-(8), and besides $\Psi({\bf r},t,q)$ can be expanded into a Taylor series with respect to $q$ at $q=0$, say,
$\Psi({\bf r},t,q)=\psi_{0}({\bf r},t)+\sum_{m=1}^{+\infty}\psi_{m}({\bf r},t)\;q^{m},$ (14)
where
$\psi_{m}({\bf r},t)=\left.\frac{1}{m!}\frac{\partial^{m}\Psi({\bf r},t,q)}{\partial q^{m}}\right|_{q=0}.$ (15)
Note that a Taylor series often has a finite radius of convergence so that the Taylor series (14) might be divergent at $q=1$. Fortunately, in the frame of the HAM, one has great freedom to choose the so-called “convergence-control parameter” $c_{0}$ that has no physical meanings. Assuming that the guess solution $\psi_{0}({\bf r},t)$, the auxiliary linear operator ${\cal L}^{\star}$ and the “convergence-control parameter” $c_{0}$ are so properly chosen that the Taylor series (14) is convergent at $q=1$, one has due to (13) the homotopy series solution
$\psi({\bf r},t)=\psi_{0}({\bf r},t)+\sum_{m=1}^{+\infty}\psi_{m}({\bf r},t).$ (16)
In practice, only finite terms are used. So, one has the $M$th-order homotopy approximation
$\psi({\bf r},t)\approx\psi_{0}({\bf r},t)+\sum_{m=1}^{M}\psi_{m}({\bf r},t).$ (17)
Note that the HAM provides us the great freedom to choose the guess solution $\psi_{0}({\bf r},t)$. Logically, the $M$th-order homotopy approximation should be a better approximation than the guess solution $\psi_{0}({\bf r},t)$, as long as the convergence-control parameter $c_{0}$ is properly chosen. So, one can use it as a new guess solution to further gain a better $M$th-order homotopy approximation. This provides us a $M$th-order iterative formula, which works quite well even at a low order [5, 6]. Here the key point is that $\psi_{m}({\bf r},t)$ in (16) is governed by a linear PDE, as described below in details.
Differentiating both sides of the zeroth-order deformation equations (9) - (11) $m$ times with respect to $q$ and then dividing them by $m!$ and finally setting $q=0$, one has the linear $m$th-order deformation equation ($m\geq 1$)
${\cal L}^{\star}\left[\psi_{m}({\bf r},t)-\chi_{m}\;\psi_{m-1}({\bf r},t)\right]=c_{0}\;\Delta_{m-1}({\bf r},t),$ (18)
subject to the initial condition
$\psi_{m}({\bf r},0)=\left\{\begin{array}[]{cl}\alpha({\bf r})-\psi_{0}({\bf r},0)&\mbox{when $m=1$}\\ 0&\mbox{otherwise}\end{array}\right.$ (19)
and the boundary condition
$\psi_{m}=\left\{\begin{array}[]{cl}\beta(t)-\psi_{0}({\bf r},t)&\mbox{when $m=1$ and ${\bf r}\in\Gamma$}\\ 0&\mbox{otherwise}\end{array}\right.$ (20)
where
$\chi_{m}=\left\{\begin{array}[]{ll}0&\mbox{when $m\leq 1$}\\ 1&\mbox{otherwise}\end{array}\right.$ (21)
and
$\Delta_{0}({\bf r},t)$ $=$ $\frac{\partial\psi_{0}({\bf r},t)}{\partial t}-{\cal N}[\psi_{0}({\bf r},t)]-g({\bf r},t),$ (22)
$\Delta_{k}({\bf r},t)$ $=$ $\frac{\partial\psi_{k}({\bf r},t)}{\partial t}-\frac{1}{k!}\frac{\partial^{k}{\cal N}[\Psi({\bf r},t,q)]}{\partial q^{k}},\hskip 21.33955ptk\geq 1.$ (23)
Note that $\Delta_{k}({\bf r},t)$ is a function dependent upon $\psi_{0}({\bf r},t),\psi_{1}({\bf r},t),\cdots,\psi_{k}({\bf r},t)$ and thus is known for the linear $m$th-order deformation equation.
It should be emphasized that the $m$th-order deformation equations (18) - (20) are linear! Since the HAM provides the great freedom to choose the auxiliary linear operator [31, 32, 30], one certainly can choose such an auxiliary linear operator
${\cal L}^{\star}[\psi({\bf r},t)]=\frac{\partial\psi({\bf r},t)}{\partial t}-{\cal L}[\psi({\bf r},t)],$ (24)
where $\cal L$ is a linear operator, which we also have great freedom to choose. Then, the linear $m$th-order deformation equation (18) can be written as
$\frac{\partial\delta_{m}({\bf r},t)}{\partial t}={\cal L}[\delta_{m}({\bf r},t)]+f_{m}({\bf r},t),$ (25)
where
$f_{m}({\bf r},t)=c_{0}\;\Delta_{m-1}({\bf r},t)$ (26)
is a known function and
$\delta_{m}({\bf r},t)=\psi_{m}({\bf r},t)-\chi_{m}\;\psi_{m-1}({\bf r},t),$ (27)
which gives
$\psi_{m}({\bf r},t)=\delta_{m}({\bf r},t)+\chi_{m}\;\psi_{m-1}({\bf r},t).$ (28)
Note that the linear PDE (25) is exactly the same as (2), and thus can be solved with quantum simulation by Schrödingerisation [1, 2, 3]. In this way, an approximate solution of the original nonlinear PDEs (6) - (8) can be obtained by quantum simulation. If this quantum simulation result is not accurate enough, one can further use it as a new guess solution $\psi_{0}({\bf r},t)$ to gain a better approximation by quantum simulation, and so on. The key point here is that the HAM can guarantee the convergence of the solution series or the iteration approach by choosing a proper value of the so-called “convergence-control parameter” $c_{0}$, as illustrated in several thousands of HAM publications [5, 6, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46].
3 Conclusions and discussions
Currently, Jin et al. proposed a quantum simulation technique for any a linear PDE, called Schrödingerisation [1, 2, 3], which has been successfully applied to solve many non-Hamiltonian linear PDEs [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. In this paper, the Schrödingerisation technique of quantum simulation is expanded to any a nonlinear PDE by means of combining the Schrödingerisation technique with the homotopy analysis method (HAM) [4, 5, 6] that can transfer any a nonlinear PDE into a series of linear PDEs with convergence guarantee of series solution. In this way, any a nonlinear PDE can be solved by quantum simulation using a quantum computer – yet to be developed in the future. For simplicity, we call it “the HAM-Schrödingerisation quantum algorithm”.
Note that it is rather time-consuming to solve PDEs related to turbulent flows by means of classical simulation algorithms such as direct numerical simulation (DNS) [47] and the clean numerical simulation (CNS) [48, 49, 50]. Hopefully, the HAM-Schrödingerisation quantum algorithm proposed in this paper can open a door to more efficiently investigate complicated turbulent flows by quantum computer someday in future.